# Leveraging Joint Predictive Embedding and Bayesian Inference in Graph Self Supervised Learning

### v2

@@ I Introduction
Graph representation learning has found widespread adoption in Social Network Analysis, Recommendation Systems, Computer Vision and Natural Language Processing[1]. It aims to learn low-dimensional embeddings of nodes or subgraphs while preserving their underlying spatial and spectral features. These learned embeddings can then be used on downstream tasks such as node classification[2], link prediction[3] and community detection[4] by training task-specific decoders or classification layers on the learned embeddings keeping the backbone frozen. Such an approach reduces the computational complexity and training time on downstream tasks. Though Graph Neural Networks(GNNs) have gained popularity over time, they require a certain number of labelled nodes to perform well. Further, owing to the complexity of graph representations spatially, it is often challenging to pre-train graphs and transfer the learned embeddings to downstream tasks.
Graph Self Supervised learning has been widely studied in the literature to facilitate graph representation learning. This includes methods such as Node2Vec, DGI[5], MVGRL[6], GRACE[7] etc. Although these methods have achieved state-of-the-art results on several social networks, they rely heavily on either graph reconstruction/feature reconstruction, re-masking or generating contrastive views by negative sampling, which is a computationally expensive process on large graphs. Graph-SSL methods that rely on reconstruction (generative methods) are decoder variant, meaning they depend on the decoder architecture(inner product decoder or symmetrical feature reconstruction). Such networks might require additional feature propagation steps such as skip connections to prevent vanishing gradients and retain learned information. Furthermore, most methods tend to use more layers or stack multiple encoders to capture long-range node interactions, which increases training complexity and leads to poor performance in downstream tasks due to over-smoothing. This diminishes the model’s ability to effectively discriminate between nodes effectively[8].
To address the limitations of existing methods, we introduce a joint embedding predictive framework that predicts subgraph embeddings of randomly sampled targets given a context conditioned on a latent variable $z$. The framework employs two encoders: a context encoder, which processes subgraphs sampled by randomly dropping nodes and updates its parameters dynamically through gradient descent, and a target encoder, whose weights are maintained as a moving average of the context encoder. The target encoder processes the original graph (without any view augmentations) to generate target node representations. Subsequently, three subgraphs are sampled from the latent space by randomly masking the target node representations. These sampled subgraphs are then fed into a global mean pooling operation to compute their embeddings, which serve as targets for a single context. To prevent representational collapse, we incorporate positional information of target nodes into the context embeddings before deriving subgraph views. A predictor network then maps the context embeddings with target positional information to the corresponding target embeddings. However, optimizing solely on context-target embeddings risks overlooking the graph’s semantic information. To overcome this, we enhance node representations by introducing an additional term to the objective function, scoring pseudo-labels predicted from the node embeddings. This is achieved by fitting a Gaussian Mixture Model (GMM) on the embeddings and evaluating the contribution of each latent feature to the pseudo-labels. Through this design, we present a graph augmentation and view invariant self-supervised learning technique that avoids the need for negative samples or contrastive objectives based on mutual information estimators. This offers a robust alternative for representation learning in graph-based applications.
@@
  • We introduce a novel graph self supervised learning method based on a joint predictive embedding paradigm which bypasses contrastive objectives such as Mutual Information(MI) Estimators, contrastive example generation techniques such as negative sampling and avoids noisy features.
  • Through our technique, we account for representation collapse by sampling multiple target embeddings for a single context, thereby enhancing the spread of node representations, and by providing positional information of target nodes to the context subgraph.
  • We also account for contribution of node embeddings to pseudo-label level predictions by incorporating an additional term to the joint predictive objective. This is given by scoring pseudo-labels generated by fitting a Gaussian Mixture Model(GMM) on the learned node embeddings.
  • We conduct extensive experiments demonstrating that our approach significantly outperforms previous state-of-the-art Graph Self-Supervised Learning (G-SSL) methods. Additionally, we provide a thorough emperical efficiency analysis highlighting the model’s efficiency and scalability.
@@ II-A Unsupervised Representation Learning on Graphs
[9] proposes a method for composing multiple self-supervised tasks for GNNs. They introduce a pseudo-homophily measure to evaluate representation quality without labels. [10] proposed the removal of negative sampling and MI estimator optimization entirely. The authors propose a loss function that contains an invariance term that maximizes correlation between embeddings of the two views and a decorrelation term that pushes different feature dimensions to be uncorrelated. [11] employs a re-mask decoding strategy and uses expressive GNN decoders instead of Multi-Layer Perceptrons(MLPs) enabling the model to learn more meaningful compressed representations. It lays focus on masked feature reconstruction rather than structural reconstruction. [6] proposed MVGRL that makes use of two graph views and a discriminator to maximize mutual information between node embeddings from a first-order view and graph embeddings from a second-order view. It leverages both node and graph-level embeddings and avoids the need for explicit negative sampling. [12] proposed ParetoGNN which simultaneously learns from multiple pretext tasks spanning different philosophies. It uses a multiple gradient descent algorithm to dynamically reconcile conflicting learning objectives, showing state-of-the-art performance in node classification, clustering, link prediction and community prediction.
@@ II-B Bootstrapping Methods
[13] introduced multi-scale feature propagation to capture long-range node interactions without oversmoothing. The authors also enhance inter-cluster separability and intra-cluster compactness by inferring cluster prototypes using a Bayesian non-parametric approach via Dirichlet Process Mixture Models(DPMMs). [14] makes use of simple graph augmentations such as random node feature and edge masking, making it easier to implement on large graphs while achieving state-of-the-art results. It leverages a cosine similarity-based objective to make the predicted target representations closer to the true representations. [15] introduced a Siamese network architecture comprising an online and target encoder with momentum-driven update steps for the target. The authors propose 2 contrastive objectives i.e cross-network and cross view-contrastiveness. The Cross-network contrastive objective incorporates negative samples to push disparate nodes away in different graph views to effectively learn topological information.
@@ II-C Joint Predictive Embedding Methods
Joint predictive embedding has been explored in the field of computer vision and audio recognition. [16] introduced I-JEPA which eliminates the need for hand-crafted augmentations and image reconstruction. They make use of a joint-embedding predictive model that predicts representations of masked image regions in an abstract feature space rather than in pixel space, allowing the model to focus on high-level semantic structures. It makes use of a Vision Transformer(ViT) with a multi-block masking strategy ensuring that the predictions retain semantic integrity. [17] extends the masked-modeling principle from vision to audio, enabling self-supervised learning on spectrograms. The key technical contribution is the introduction of a curriculum masking strategy, which transitions from random block masking to time-frequency-aware masking, addressing the strong correlations in audio spectrograms.
@@ III Methodology
Figure 1: Overall Framework Diagram
@@ III-B1 Joint Predictive Optimization
We sample a subgraph $G^{\prime}$ from graph $G$ by dropping a set of nodes according to a Bernoulli distribution parameterized by success probability $p_{1}$. The subgraph $G^{\prime}$ is passed into the context encoder to output node embeddings $H^{\prime}$ of dimensions $d^{\prime}$. Meanwhile, the graph $G$ is passed into the target encoder to generate node embeddings $H$ with the same dimensions as $H^{\prime}$. At the latent space, we sample 3 target subgraphs from the context subgraph $G^{\prime}$, again according to a Bernoulli distribution with probability $p_{2}$ such that $p_{2}<p_{1}$. The node embeddings of each subgraph are obtained by deactivating the node features of masked nodes, which is followed by global mean pooling operation, thus obtaining embeddings $H^{t}_{1}$,$H^{t}_{2}$ and $H^{t}_{3}$. The context node level embeddings $H^{\prime}$ is passed into 3 predictors, corresponding to the three target subgraphs. The obtained predictions are then pooled by the mean pooling operation, thus resulting in target predictions $H^{\prime}_{1},H^{\prime}_{2},H^{\prime}_{3}$, conditioned on latent variable $H^{\prime}$. We then frame the objective function for the joint predictive component as follows,
@@ III-D Implementation Details
For all experiments, we use the Adam Optimizer[18] and cosine annealing learning rate scheduler with warm restarts. No regularization techniques such as dropout, $L_{1}$ or $L_{2}$ have been employed since they tend to reduce performance. For learning rates, we perform a search with values corresponding to the search space $\{0.5,0.1,0.05,0.01,0.001\}$. We set the value of the momentum update parameter $m$ to 0.9. The number of epochs was set to 50,000 with early stopping. All experiments were conducted on an M2 Macbook Pro with 8GB CPU. We did not utilize GPUs for any of our experiments.
@@ IV-A1 Datasets
In our experiments, we evaluate our proposed framework on seven publicly available benchmark datasets for node representation learning and classification. The datasets are, namely, Cora[19], Pubmed[20], Citeseer[19], Amazon Photos[21], Amazon Computers[21], Coauthor CS[21] and WikiCS[22]. Details on number of nodes, edges and features are given in table I Cora consists of 2708 scientific publications classified into one of seven classes, Citeseer consists of 3312 scientific publications classified into one of six classes and Pubmed consists of 19717 scientific publications pertaining to diabetes classified into one of three classes. In Amazon Photos and Amazon Computers, nodes represent goods and edges represent that two goods are frequently bought together. The product reviews are represented as bag-of-words features. Coauthor CS contains paper keywords for each author’s papers. Nodes represent authors that are connected by an edge if they co-authored a paper. WikiCS is a dataset derived from Wikipedia with nodes corresponding to Computer Science articles and edges based on hyperlinks. The 10 classes represent different branches of Computer Science.
@@ IV-A2 Evaluation Methodology
The embeddings are evaluated on node classification, whose performance is quantified by accuracy. We follow the same evaluation protocol as used in [23], [13], [12] etc. All scores for baselines have been obtained from previously published papers.
@@ IV-A3 Evaluation Protocol
For all downstream tests, we follow the linear evaluation protocol on graphs where the parameters of the backbone/encoder are frozen during inference time. Only the prediction head, which is a single GCN layer, is trained for node classification. For evaluation purposes, we use the default number for train/val/test splits for the citation networks i.e. Cora, Pubmed, Citeseer which are 500 validation and 1000 testing nodes. The train/val/test splits for the remaining datasets, namely, Amazon Photos, Amazon Computers and Coauthor CS are according to [21]. For WikiCS we use the publicily available splits. Unless otherwise mentioned, performance is averaged over 10 independent runs with different random seeds and splits for all seven evaluation datasets. We report the mean and standard deviation obtained across 10 runs.
@@ IV-B1 Performance on Node Classification on small and large graphs
As mentioned earlier, we use the linear evaluation protocol and report the mean and standard deviation of classification accuracy on the test nodes over 10 runs on different folds or splits with different seeds. We compare our proposed approach with supervised and fine-tuned model baselines. For semi-supervised node classification, we compare our proposed approach against Multi Layer Perceptron(MLP),Graph Convolution Network(GCN)[24], Graph Attention Network(GAT)[23], Simplified GCN[25], Logistic Regression and GraphSAGE[26]. For GraphSAGE, we use the mean, maxpool and meanpool variants as described in [21]. We have compared our model’s performance against semi-supervised baselines in table II. For self-supervised and fine-tuned node classification, we compare our proposed approach against DGI, MVGRL[6], GRACE[7], CCA-SSG[10], SUGRL[27], S3-CL[13], GraphMAE[11], GMI[28], BGRL[14] and ParetoGNN[12] on small and large graphs. In semi-supervised node classification (Table II), our method achieves the highest accuracy across all baselines. On self-supervised classification followed by fine-tuning, we evaluate our proposed model against two specific fields of data i.e Planetoid datasets and larger, more stable datasets such as Amazon Computers, Photos and WikiCS. Table III compares our model’s performance against strong baselines on Planetoid datasets. We show that our model performs extremely well on smaller datasets by outperforming all baselines on Cora and Citeseer with extremely competitive results on Pubmed. Table IV contains the performance of our model against strong baselines for Photos, Computers, Coauthor CS and WikiCS. Our model consistently achieves competitive scores against baselines by outperforming baselines on Photos and CS while remaining competitive on Computers and WikiCS .These results underscore the approach’s adaptability and strong performance across diverse dataset sizes. Owing to the ability of our proposed approach to avoid noisy features, representation collapse, and leverage semantic information, our proposed model performs well on small and unstable datasets such as Cora and Citeseer and large datasets such as Amazon, Coauthor and WikiCS.
@@
TABLE IV: Self Supervised Pre-Training followed by classification on datasets mentioned in [21] and WikiCS. The best score for each dataset is marked in bold and the second-best score is underlined.

### v4

# Predict, Cluster, Refine: A Joint Embedding Predictive Self-Supervised Framework for Graph Representation Learning

@@ I Introduction
Graph representation learning has become a cornerstone of modern machine learning, powering advancements in domains as diverse as Social Network Analysis, Recommendation Systems, and Drug Discovery (wu2020comprehensive, ). The core promise is to distill complex, high-dimensional graph structures into potent, low-dimensional embeddings that capture essential spatial and spectral features. These embeddings can then fuel downstream tasks like node classification (maekawa2022beyond, ) and link prediction (zhang2018link, ) with remarkable efficiency, often by training simple decoders on a fixed, pre-trained backbone. However, the dominant paradigm, Graph Neural Networks (GNNs), faces a critical bottleneck: their performance is fundamentally tied to the availability of large, labeled datasets, which are often scarce and expensive to acquire. This dependency severely limits their applicability in real-world, data-sparse scenarios.
To break free from this supervised learning constraint, Graph Self-Supervised Learning (G-SSL) has emerged as a powerful alternative. Pioneers like Node2Vec, DGI (velivckovic2018deep, ), and GRACE (zhu2020deep, ) have demonstrated impressive results by learning from the graph’s intrinsic structure. Yet, a closer inspection reveals that the current G-SSL landscape is dominated by two problematic paradigms. The first, contrastive learning, relies on manually-defined graph augmentations and the generation of negative samples to pull similar nodes together and push dissimilar ones apart. This process is not only computationally prohibitive on large-scale graphs but is also highly sensitive to the choice of views and negative sampling strategy, often introducing noise and instability. The second, generative learning, focuses on reconstructing masked features or graph structures. These methods are notoriously fragile, often depending on specific decoder architectures and requiring complex designs like skip connections to combat vanishing gradients. Furthermore, both paradigms frequently resort to deep, multi-layer encoders to capture long-range dependencies, inadvertently falling prey to the pervasive issue of over-smoothing, which erodes node distinguishability and cripples downstream performance (WANG2024106484, ).
To transcend these limitations, we introduce a fundamentally different approach: a joint embedding predictive framework that sidesteps the pitfalls of both contrastive and generative methods. Our model learns by predicting the latent representations of a target subgraph from the representation of a context subgraph, conditioned on a shared latent variable. This is achieved through an elegant asymmetric architecture: a context encoder, trained via standard gradient descent on augmented graph views, and a target encoder, which generates stable, high-quality representations from the original graph and is updated as an exponential moving average of the context encoder. This design completely obviates the need for negative sampling and fragile reconstruction objectives
Crucially, our framework incorporates two novel mechanisms to ensure rich and well-distributed embeddings. First, to combat representational collapse, we predict multiple, randomly masked target subgraph embeddings from a single context, forcing the model to learn a more diverse and informative latent space. This is further enhanced by injecting the positional information of target nodes into the context. Second, to ensure the embeddings are not just discriminative but also semantically coherent, we introduce a powerful regularization term. We fit a Gaussian Mixture Model (GMM) on the learned embeddings to generate pseudo-labels and then score the contribution of each latent feature to these clusters. This forces the model to capture high-level community structure within the graph. The result is a robust, augmentation-invariant G-SSL technique that learns powerful representations without the computational overhead and design complexities of its predecessors.
@@
  • We propose a novel joint embedding predictive framework that obviates the need for contrastive objectives, negative sampling, and explicit graph reconstruction, offering a more efficient and direct learning paradigm.
  • We introduce a multi-target prediction strategy coupled with positional encoding to explicitly combat representational collapse, ensuring a well-distributed and informative embedding space.
  • We pioneer a semantic regularization objective that scores the contribution of learned embeddings to GMM-derived pseudo-labels, guiding the model to capture meaningful high-level community structures.
  • We conduct extensive experiments demonstrating that our approach significantly outperforms previous state-of-the-art G-SSL methods and provide a thorough empirical analysis highlighting its superior efficiency and scalability.
@@ II-A Unsupervised Representation Learning on Graphs
(jin2021automated, ) proposes a method for composing multiple self-supervised tasks for GNNs. They introduce a pseudo-homophily measure to evaluate representation quality without labels. (zhang2021canonical, ) proposed the removal of negative sampling and MI estimator optimization entirely. The authors propose a loss function that contains an invariance term that maximizes correlation between embeddings of the two views and a decorrelation term that pushes different feature dimensions to be uncorrelated. (hou2022graphmae, ) employs a re-mask decoding strategy and uses expressive GNN decoders instead of Multi-Layer Perceptrons(MLPs) enabling the model to learn more meaningful compressed representations. It lays focus on masked feature reconstruction rather than structural reconstruction. (hassani2020contrastive, ) proposed MVGRL that makes use of two graph views and a discriminator to maximize mutual information between node embeddings from a first-order view and graph embeddings from a second-order view. It leverages both node and graph-level embeddings and avoids the need for explicit negative sampling. (ju2022multi, ) proposed ParetoGNN which simultaneously learns from multiple pretext tasks spanning different philosophies. It uses a multiple gradient descent algorithm to dynamically reconcile conflicting learning objectives, showing state-of-the-art performance in node classification, clustering, link prediction and community prediction.
@@ II-B Bootstrapping Methods
(ding2023eliciting, ) introduced multi-scale feature propagation to capture long-range node interactions without oversmoothing. The authors also enhance inter-cluster separability and intra-cluster compactness by inferring cluster prototypes using a Bayesian non-parametric approach via Dirichlet Process Mixture Models(DPMMs). (thakoor2021large, ) makes use of simple graph augmentations such as random node feature and edge masking, making it easier to implement on large graphs while achieving state-of-the-art results. It leverages a cosine similarity-based objective to make the predicted target representations closer to the true representations. (jin2021multi, ) introduced a Siamese network architecture comprising an online and target encoder with momentum-driven update steps for the target. The authors propose 2 contrastive objectives i.e cross-network and cross view-contrastiveness. The Cross-network contrastive objective incorporates negative samples to push disparate nodes away in different graph views to effectively learn topological information.
@@ II-C Joint Predictive Embedding Methods
Joint predictive embedding has been explored in the field of computer vision and audio recognition. (assran2023self, ) introduced I-JEPA which eliminates the need for hand-crafted augmentations and image reconstruction. They make use of a joint-embedding predictive model that predicts representations of masked image regions in an abstract feature space rather than in pixel space, allowing the model to focus on high-level semantic structures. It makes use of a Vision Transformer(ViT) with a multi-block masking strategy ensuring that the predictions retain semantic integrity. (fei2023jepa, ) extends the masked-modeling principle from vision to audio, enabling self-supervised learning on spectrograms. The key technical contribution is the introduction of a curriculum masking strategy, which transitions from random block masking to time-frequency-aware masking, addressing the strong correlations in audio spectrograms.
@@ III Methodology
Figure 1: The Proposed Self-Supervised Learning Architecture. The framework operates in two parallel streams. The primary predictive task involves an asymmetric encoder design. A Context Encoder processes an augmented graph view to produce a context embedding. This is used to predict the representations of multiple, distinct Target Subgraphs, which are generated by a momentum-based Target Encoder from the original graph. This multi-target prediction prevents the model from learning trivial solutions. A Semantic Consistency Module runs in parallel, fitting a Gaussian Mixture Model (GMM) on the learned node embeddings to derive pseudo-labels. These labels are then used to compute a regularization term that re-weights the embeddings, ensuring they capture coherent, high-level graph structures. The final objective is a weighted sum of the predictive loss and the semantic consistency score.
@@ III-B1 Joint Predictive Optimization
We sample a subgraph $G^{\prime}$ from graph $G$ by dropping a set of nodes according to a Bernoulli distribution parameterized by success probability $p_{1}$. The subgraph $G^{\prime}$ is passed into the context encoder to output node embeddings $H^{\prime}$ of dimensions $d^{\prime}$. Meanwhile, the graph $G$ is passed into the target encoder to generate node embeddings $H$ with the same dimensions as $H^{\prime}$. At the latent space, we sample 3 target subgraphs from the context subgraph $G^{\prime}$, again according to a Bernoulli distribution with probability $p_{2}$ such that $p_{2}<p_{1}$. The node embeddings of each subgraph are obtained by deactivating the node features of masked nodes, which is followed by global mean pooling operation, thus obtaining embeddings $H^{t}_{1}$$H^{t}_{2}$ and $H^{t}_{3}$. The context node level embeddings $H^{\prime}$ is passed into 3 predictors, corresponding to the three target subgraphs. The obtained predictions are then pooled by the mean pooling operation, thus resulting in target predictions $H^{\prime}_{1},H^{\prime}_{2},H^{\prime}_{3}$, conditioned on latent variable $H^{\prime}$. We then frame the objective function for the joint predictive component as follows,
@@ III-D Implementation Details
For all experiments, we use the Adam Optimizer(kingma2014adam, ) and cosine annealing learning rate scheduler with warm restarts. No regularization techniques such as dropout, $L_{1}$ or $L_{2}$ have been employed since they tend to reduce performance. For learning rates, we perform a search with values corresponding to the search space $\{0.5,0.1,0.05,0.01,0.001\}$. We set the value of the momentum update parameter $m$ to 0.9. The number of epochs was set to 50,000 with early stopping. All experiments were conducted on an M2 Macbook Pro with 8GB CPU.
@@ IV-A1 Datasets
In our experiments, we evaluate our proposed framework on seven publicly available benchmark datasets for node representation learning and classification. The datasets are, namely, Cora(sen2008collective, ), Pubmed(namata2012query, ), Citeseer(sen2008collective, ), Amazon Photos(shchur2018pitfalls, ), Amazon Computers(shchur2018pitfalls, ), Coauthor CS(shchur2018pitfalls, ) and WikiCS(mernyei2020wiki, ). Details on number of nodes, edges and features are given in table I. Cora consists of 2708 scientific publications classified into one of seven classes, Citeseer consists of 3312 scientific publications classified into one of six classes and Pubmed consists of 19717 scientific publications pertaining to diabetes classified into one of three classes. In Amazon Photos and Amazon Computers, nodes represent goods and edges represent that two goods are frequently bought together. The product reviews are represented as bag-of-words features. Coauthor CS contains paper keywords for each author’s papers. Nodes represent authors that are connected by an edge if they co-authored a paper. WikiCS is a dataset derived from Wikipedia with nodes corresponding to Computer Science articles and edges based on hyperlinks. The 10 classes represent different branches of Computer Science.
@@ IV-A2 Evaluation Methodology
The embeddings are evaluated on node classification, whose performance is quantified by accuracy. We follow the same evaluation protocol as used in (velivckovic2017graph, ), (ding2023eliciting, ), (ju2022multi, ) etc. All scores for baselines have been obtained from previously published papers.
@@ IV-A3 Evaluation Protocol
For all downstream tests, we follow the linear evaluation protocol on graphs where the parameters of the backbone/encoder are frozen during inference time. Only the prediction head, which is a single GCN layer, is trained for node classification. For evaluation purposes, we use the default number for train/val/test splits for the citation networks i.e. Cora, Pubmed, Citeseer which are 500 validation and 1000 testing nodes. The train/val/test splits for the remaining datasets, namely, Amazon Photos, Amazon Computers and Coauthor CS are according to (shchur2018pitfalls, ). For WikiCS we use the publicly available splits. Unless otherwise mentioned, performance is averaged over 10 independent runs with different random seeds and splits for all seven evaluation datasets. We report the mean and standard deviation obtained across 10 runs.
@@ IV-B1 Performance on Node Classification on small and large graphs
As mentioned earlier, we use the linear evaluation protocol and report the mean and standard deviation of classification accuracy on the test nodes over 10 runs on different folds or splits with different seeds. We compare our proposed approach with supervised and fine-tuned model baselines. For semi-supervised node classification, we compare our proposed approach against Multi Layer Perceptron(MLP),Graph Convolution Network(GCN)(kipf2016semi, ), Graph Attention Network(GAT)(velivckovic2017graph, ), Simplified GCN(wu2019simplifying, ), Logistic Regression and GraphSAGE(hamilton2017inductive, ). For GraphSAGE, we use the mean, maxpool and meanpool variants as described in (shchur2018pitfalls, ). We have compared our model’s performance against semi-supervised baselines in table II. For self-supervised and fine-tuned node classification, we compare our proposed approach against DGI, MVGRL(hassani2020contrastive, ), GRACE(zhu2020deep, ), CCA-SSG(zhang2021canonical, ), SUGRL(mo2022simple, ), S3-CL(ding2023eliciting, ), GraphMAE(hou2022graphmae, ), GMI(peng2020graph, ), BGRL(thakoor2021large, ) and ParetoGNN(ju2022multi, ) on small and large graphs. In semi-supervised node classification (Table II), our method achieves the highest accuracy across all baselines. On self-supervised classification followed by fine-tuning, we evaluate our proposed model against two specific fields of data i.e Planetoid datasets and larger, more stable datasets such as Amazon Computers, Photos and WikiCS. Table III compares our model’s performance against strong baselines on Planetoid datasets. We show that our model performs extremely well on smaller datasets by outperforming all baselines on Cora and Citeseer with extremely competitive results on Pubmed. Table IV contains the performance of our model against strong baselines for Photos, Computers, Coauthor CS and WikiCS. Our model consistently achieves competitive scores against baselines by outperforming baselines on Photos and CS while remaining competitive on Computers and WikiCS .These results underscore the approach’s adaptability and strong performance across diverse dataset sizes. Owing to the ability of our proposed approach to avoid noisy features, representation collapse, and leverage semantic information, our proposed model performs well on small and unstable datasets such as Cora and Citeseer and large datasets such as Amazon, Coauthor and WikiCS.
@@
TABLE IV: Self Supervised Pre-Training followed by classification on datasets mentioned in (shchur2018pitfalls, ) and WikiCS. The best score for each dataset is marked in bold and the second-best score is underlined.